1. Jim drops a stone from the top of a five-storey building (a) How far has it fallen after 2 seconds? (b) What is the final velocity of the stone after 2 seconds of fall? (c) if the stone reaches the ground after 3 seconds, how high is the building? It is worth at this point to make a simple sketch.


2.Just follow the instruction.

You may try this with your siblings. You only need 1 meter string, a thread or nylon cord.


1. Ask for volunteers


2. One learner would hold a nylon cord at length across two hands.


3. The second learner loops his nylon cord onto the other learner's cord.


4. The second learner pulls slowly on the cord, if the loop is closer to the other learner's hand, how would you feel the pull on each hand, and why? Hint: Stronger force, in this case the tension, must be represented by longer vector.


1
Expert's answer
2021-09-10T12:18:36-0400

1. The distance is


"h=\frac{gt^2}{2}=19.6\text{ m}."

The final velocity is


"v=\sqrt{2gh}=19.6\text{ m/s}."

The height of the building is


"H=\frac{gT^2}{2}=44.1\text{ m}."

2. I would feel the pull differently on both hands: stronger pull on the hand which is closer to the place where the second cord is looped onto mine.